For the pair of vectors, find 3U-4V U=(3,2), V=(-4,3) a. (25, -6) ob. (25.6) Oc(24,-5) O d. (-24,5) O e. (6. – 25)
Answer #1

U= (3,2) (-4,3) v 30-4V = 3 (3,2) - 4 64,3) (9,6)+ (16,42) - (9+16, 6-12) =<25,-67 30-4V 528,-6) Therefore option option @ is